OpenStudy (epicteatime):

What is the simplified form of the expression sqrt(16c^8) - (c^2) + 5 sqrt(c~2) + (4c^2)

OpenStudy (epicteatime):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Simplify the expressions in the square root, first.

OpenStudy (epicteatime):

ok ;3

OpenStudy (ipwnbunnies):

\[\sqrt{16c^{8}} = \sqrt{4*4*c^{4}*c^{4}} = 4c^{4}\]

OpenStudy (epicteatime):

ok i got it

OpenStudy (ipwnbunnies):

\[5 \sqrt{c^{2}} = 5c\] Now, combine like terms. c:

OpenStudy (epicteatime):

do i just add them all together?

OpenStudy (ipwnbunnies):

Yes, that's what combine like terms are. :3 But remember, you can only add variables of a similar power. So, you can't add c^4's and c^2's together.

OpenStudy (epicteatime):

ok ;L

OpenStudy (ipwnbunnies):

Now we have: \[4c^{4} - c^{2} + 5c + 4c^{2}\]

OpenStudy (ipwnbunnies):

Add similar terms. The c^2 terms will combine, in this case. :3

OpenStudy (epicteatime):

yes

OpenStudy (epicteatime):

4c^2+5c+3c^2?

OpenStudy (ipwnbunnies):

Yesh. c:

OpenStudy (epicteatime):

oh

OpenStudy (epicteatime):

so thats the answer?

OpenStudy (ipwnbunnies):

That is the simplified form of the expression. o-o

OpenStudy (epicteatime):

hmmm

OpenStudy (ipwnbunnies):

Unless you can factor out a whole number, which you can't. The coefficents of the variables don't share a common factor.

OpenStudy (epicteatime):

A) 8c4 + 8c B) 4c4 + 8c C) 8c4 + 6c D) 4c4 + 6 ;P;P;P;P;P;P;P;P;P

OpenStudy (epicteatime):

hmmmmmmm

OpenStudy (epicteatime):

from my opinion its either c or d

OpenStudy (epicteatime):

cause thats the only logical answer

OpenStudy (ipwnbunnies):

Did you write the problem out correctly?

OpenStudy (epicteatime):

i have nooo idea

OpenStudy (ipwnbunnies):

o-o

OpenStudy (epicteatime):

\[\sqrt{16c ^{8}}-\sqrt{c ^{2}}+5\sqrt{c ^{2}}+\sqrt{4c ^{2}}\]

OpenStudy (epicteatime):

there ;P

OpenStudy (epicteatime):

is that better? or does it make it more confusing?

OpenStudy (ipwnbunnies):

Ahh, there we go. ^.^ Again, simplify the square roots. Eliminate them. :3

OpenStudy (epicteatime):

ooo

OpenStudy (epicteatime):

so does it become the same answer?

OpenStudy (ipwnbunnies):

No, you left out a few square roots in the problem you originally posted.

OpenStudy (epicteatime):

oh........

OpenStudy (ipwnbunnies):

Can you not do the work? Simplify square roots? o-o

OpenStudy (epicteatime):

NUUUUUUUUUUUUUUUUU~~~~~~!!!!!!!!! 0O0'

OpenStudy (ipwnbunnies):

\[\sqrt{c^2} = \sqrt{c*c} = c\] There are two c's being multiplied under a square root. You can pull them both out, and it just becomes c.

OpenStudy (epicteatime):

4c^4

OpenStudy (epicteatime):

yeah i know that one

OpenStudy (ipwnbunnies):

Mhm, go on. :3

OpenStudy (epicteatime):

problem is can you add the sqrt(4c^2) with the other c^2 thing

OpenStudy (epicteatime):

you can't right?

OpenStudy (epicteatime):

and then the last one is 2c right?

OpenStudy (ipwnbunnies):

You can only add square roots when the thing under the radical is the same. For example: \[5\sqrt{x^{2}} + 2\sqrt{x^{2}} = 7\sqrt{x^{2}}\]

OpenStudy (ipwnbunnies):

Yes, last one is 2c.. c: NOW, combine like terms again. It'll be a nice answer.

OpenStudy (epicteatime):

yay!

OpenStudy (epicteatime):

answers op!

OpenStudy (epicteatime):

thanks i got it ;P so it was D

OpenStudy (epicteatime):

my prediction was correct, i am better than SH!

OpenStudy (epicteatime):

MWAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHAHA!

OpenStudy (ipwnbunnies):

o_o

OpenStudy (epicteatime):

can i also ask you another question tho?

OpenStudy (ipwnbunnies):

Sure.

OpenStudy (epicteatime):

\[2\sqrt{10k}\times5k \sqrt{2}\]

OpenStudy (epicteatime):

there ;3

OpenStudy (epicteatime):

hmmmmmm

OpenStudy (ipwnbunnies):

When multiplying radicals, multiply the stuff outside the radicals. Then, multiply the stuff inside the radicals, and put that product under another radical.

OpenStudy (epicteatime):

ok

OpenStudy (ipwnbunnies):

And then you can simplify what's under the radical even more, in this problem.

OpenStudy (epicteatime):

but

OpenStudy (epicteatime):

how do i multiply whats inside the radicals

OpenStudy (epicteatime):

do i multiply 10k and2?

OpenStudy (ipwnbunnies):

10k is one factor. 2 is the other factor. o-o Yes

OpenStudy (epicteatime):

so i X it together?

OpenStudy (ipwnbunnies):

Yesh.

OpenStudy (epicteatime):

which makes 20k

OpenStudy (epicteatime):

then 10k?

OpenStudy (epicteatime):

which is \[10k \sqrt{20k}\]

OpenStudy (ipwnbunnies):

Correct. c: Can you simplify the square root now?

OpenStudy (epicteatime):

hmmm

OpenStudy (epicteatime):

2k?

OpenStudy (epicteatime):

tooshay?

OpenStudy (epicteatime):

;3

OpenStudy (ipwnbunnies):

Nope. :3 You can't pull out a 'k' from the square root if there's only 1 'k' under the square root.

OpenStudy (ipwnbunnies):

\[\sqrt{20k} = \sqrt{4*5*k} = \sqrt{2*2*5*k}\]

OpenStudy (epicteatime):

20k sqrt5

OpenStudy (epicteatime):

5k

OpenStudy (ipwnbunnies):

Yesh. ^-^

OpenStudy (epicteatime):

hooray!

OpenStudy (epicteatime):

\[\sqrt{2}(6b \sqrt{t}-5)\]

OpenStudy (epicteatime):

oops not 6b

OpenStudy (ipwnbunnies):

Distribute the sqrt(2) to each term.

OpenStudy (epicteatime):

sqrt(2) (8sqrt(t)-5)

OpenStudy (epicteatime):

\[8\sqrt{2t}-5\sqrt{2}\]

OpenStudy (ipwnbunnies):

You got it. ^--^

OpenStudy (epicteatime):

yay!

OpenStudy (epicteatime):

do i add sqrt's together?

OpenStudy (anonymous):

Wow all of this instead of going here. http://bit.ly/1q1JkSH

OpenStudy (epicteatime):

like sqrt(2) and sqrt(16)

OpenStudy (epicteatime):

so

OpenStudy (epicteatime):

do i add the sr together? or multiply them?

OpenStudy (ipwnbunnies):

You can only add them if the stuff under the radical is the same.

OpenStudy (epicteatime):

so its 18?

OpenStudy (ipwnbunnies):

Nooo. You can't add those together. Only if terms SHARE the same square root 'stuff', you can add them. I showed you earlier. \[5\sqrt{2x} + 4\sqrt{x} + 2\sqrt{2x} = 6\sqrt{2x} + 4\sqrt{x}\] As an exmaple

OpenStudy (ipwnbunnies):

Omg. 5+2 = 7, not 6. ._.

OpenStudy (epicteatime):

OH NO

OpenStudy (epicteatime):

ThE MATHLETE MADE A MISTAKE THE WORLDS GONNA DIE!

OpenStudy (epicteatime):

AHHHHWUDFUSVJGEFHUIWKSCJBHJFEWKYIR#$'"()EQODUISVHGYU#R"E